OpenStudy (anonymous):

What is the probability of drawing yellow marble followed by a red marble from a bag containing 12 yellow marbles, 14 red marbles, and 15 green marbles if the first marble is not replaced? 18/43 21/205 192/1,849

OpenStudy (anonymous):

I got 18/43..?

OpenStudy (cwrw238):

total number of marbles = 12 +14+15 = 41 as there are 12 marbles what do you think the probability of drawing a yellow one is

OpenStudy (cwrw238):

18/43 is not correct

OpenStudy (cwrw238):

the probability of frist drawing a yellow is 12/41 be cause there are 12 yellows in a total of 41 marbles do you follow that?

OpenStudy (anonymous):

But there is no answer choice for that

OpenStudy (cwrw238):

i know that is on;y the first part of the answer - but you do follow that right?

OpenStudy (anonymous):

Yes i understand

OpenStudy (cwrw238):

ok - so thats not replaced so there is now 40 marbles for the second draw and there are 14 red marbles so in a similar way to to drawing the yellow one , can you work out the probability of drawing a red marble?

OpenStudy (anonymous):

just 14/40?

OpenStudy (cwrw238):

for the yellow we had 12 yellows and total 41 - probability 12/41 now for the red we have 14 reds and total of 40 - probability of ?

OpenStudy (cwrw238):

yes right 14/40

OpenStudy (cwrw238):

that the second half of the problem

OpenStudy (cwrw238):

There is a law of probibilty which states that for 2or more independent events the probabilty of the 2 occurring are MULTIPLIED the 2 events here are the yellow draw then the red and they are independent of each other so probability( yellow then red ) = 12/41 * 14/40 which if you use ur calculator you'll find it comes to 21 / 205

OpenStudy (anonymous):

Okay I see, Thank you for explaining! much simpler than it looks!

OpenStudy (cwrw238):

it is